# Dissipating quartets of excitations in a superconducting circuit

### v2

@@ Acknowledgments
Supplemental Materials
S1 Circuit design and measurement setup
S1.1 Target parameters and circuit design
@@
${\bf H}_{4}=\hbar g_{4}({\bf a}^{4}-\alpha^{4}){\bf b^{\dagger}}+\mathrm{h.c.}$ (S1)
@@
$\dot{\bm{\rho}}=\gamma_{4}\mathcal{D}[{\bf a}^{4}-\alpha^{4}]\bm{\rho}$ (S2)
with $\gamma_{4}=4g_{4}^{2}/\kappa_{b}$, and the coherent states spanning the cat manifold are attractors of the dynamics. Note that this property remains true in the regime $g_{4}\gtrsim\kappa_{b}$, even though the two-mode dynamics is no longer accurately captured by (S2[31]. In Eq.(S1), the term $\alpha^{4}{\bf b^{\dagger}}+\mathrm{h.c.}$ is trivially implemented with a resonant drive on the buffer, so that we now focus on the engineering of the more challenging four-to-one photon exchange interaction.
@@
${\bf H}(t)=\hbar\omega_{a}{\bf a^{\dagger}}{\bf a}+\hbar\omega_{b}{\bf b}{\bf b^{\dagger}}+2E_{J}\epsilon(t)\sin(\varphi^{ZPF}_{a}({\bf a}+{\bf a^{\dagger}})+\varphi^{ZPF}_{b}({\bf b}+{\bf b^{\dagger}}))$ (S3)
where we have neglected the weakly participating readout mode and considered an ideal circuit for simplicity (see Sec. S2 for an analysis of non-idealities). Note that this expression is valid in the limit $|\epsilon(t)|\ll 1$ as it results from a linearization of $\sin(\epsilon(t))$ [25].
@@
$\begin{split}e^{i\varphi^{ZPF}_{a}({\bf a}e^{-i\omega_{a}t}+{\bf a^{\dagger}}e^{+i\omega_{a}t})}&=e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}}\sum_{n,m}\frac{(i{\varphi_{a}^{ZPF}})^{n+m}}{n!m!}e^{i(n-m)\omega t}{\bf a^{\dagger}}^{n}{\bf a}^{m}\\ e^{-i{\varphi_{a}^{ZPF}}({\bf a}e^{-i\omega_{a}t}+{\bf a^{\dagger}}e^{+i\omega_{a}t})}&=e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}}\sum_{n,m}\frac{(-i{\varphi_{a}^{ZPF}})^{n+m}}{n!m!}e^{i(n-m)\omega t}{\bf a^{\dagger}}^{n}{\bf a}^{m}\end{split}$ (S4)
@@
$\tilde{{\bf H}}_{4}={\hbar}g_{4}:P_{a}({\bf a^{\dagger}}{\bf a})P_{b}({\bf b^{\dagger}}{\bf b})~{}{\bf a}^{4}{\bf b^{\dagger}}:~{}+~{}\mathrm{h.c.}$ (S5)
where :_::\_:: _ : denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi{\varphi_{a}^{ZPF}}^{4}{\varphi_{b}^{ZPF}}e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}-\frac{{\varphi_{b}^{ZPF}}^{2}}{2}}$ and
$\begin{split}P_{a}(X)={4!}\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{a}^{ZPF}}^{2n}}{n!(n+4)!}X^{n}\\ P_{b}(X)=\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{b}^{ZPF}}^{2n}}{n!(n+1)!}X^{n}\end{split}$ (S6)
Compared to the target Hamiltonian (S1), higher-order terms appear (corresponding to terms indexed by $n\geqslant 1$ in $P_{a}$ and $P_{b}$). These terms renormalize the four-to-one photon exchange rate between states $|k+4\rangle_{a}|l\rangle_{b}$ and $|k\rangle_{a}|l+1\rangle_{b}$ for $(k,l)\neq(0,0)$. Note that in our experiment, the modes are non-linear so that the pump is not resonant with any of these transitions. Therefore, the system dynamics is the same as that yielded by (S1) (see Sec. S2 and Sec. S4.2).
The four-to-one photon exchange rate $g_{4}$ in (S5) depends strongly on ${\varphi_{a}^{ZPF}}$, which should thus be maximized. However, when stabilizing a four-legged cat in a linear system, high-order terms in the sine expansion can have a a significant impact on the dynamics when $\varphi^{ZPF}_{a}\alpha\gtrsim 1$ such that the sine potential cannot be truncated at leading order. In practice, we saw in numerical simulations that beyond this limit, the cat blobs are distorted and eventually become unstable (not shown). In particular, we found that one should limit ${\varphi_{a}^{ZPF}}\leq 0.5$ to stabilize a cat containing up to 10 photons on average. As for high-order terms in $P_{b}$, they should have no impact on the system dynamics when it is close to its steady state [72]. However, the prefactor $e^{-\frac{{\varphi_{b}^{ZPF}}^{2}}{2}}$ limits the four-to-one photon exchange rate which reaches a maximum for $\varphi^{ZPF}_{b}=1$. In practice, designing a circuit with modes whose vacuum fluctuations across the ATS approach 1 is challenging and one gains only marginally beyond $\varphi^{ZPF}_{b}\sim 0.5$, so that we targeted in design ${\varphi_{a}^{ZPF}}=0.45$ and ${\varphi_{b}^{ZPF}}=0.55$.
As for the readout mode which we have neglected so far, we targeted ${\varphi_{r}^{ZPF}}=0.05$ which was deemed sufficient to perform photon-number resolved measurements of the memory leveraging the non-linearity of the ATS  [^8] We considered activating a longitudinal coupling between the memory and buffer by pumping the ATS at $\omega_{b}$, following the strategy presented in Ref. [27]. In practice this measurement scheme proved hard to tune in our system with significant dispersive interaction between modes, and dispersive measurements were sufficient to achieve the same goal while ensuring that the memory anharmonicity remained smaller than its linewidth at all the ATS DC bias points. This last property is convenient to characterize the system by monitoring the readout mode frequency as a function of bias fluxes (see Fig. 2).
@@
The mode resonance frequencies are not particularly constrained toward the goal of stabilizing a four-legged cat. However, the resonance frequencies should be sufficiently separated that the feedlines can be filtered to obtain a strong single-photon dissipation on the buffer while retaining a large intrinsic quality factor for the memory. Moreover, we generate the pump microwave with standard microwave equipment (in the $1-20$ GHz range) and should avoid frequency collisions with processes resonant at $n\omega_{a}+m\omega_{b}$ with $(n,m)\neq(4,-1)$. In practice, we targeted $\omega_{a}=2\pi\times 4.5~{}$GHz and $\omega_{b}=2\pi\times 8~{}$GHz. For the resonance frequency of the readout mode, we aimed at the center of the working range of our parametric amplifier (see Fig. S3).
@@
The four-to-one photon exchange rate $g_{4}$ scales linearly with $E_{J}$ so that naively, one would aim for large $E_{J}$ values to increase the engineered dissipation rate. However, spurious non-linearities stemming from the tunneling of pairs of Cooper pairs through the SQUID scales with $E_{J}^{2}$ (see Sec. S2). Through numerical simulations, we estimated that for the target values of ${\varphi_{a}^{ZPF}}$ and ${\varphi_{b}^{ZPF}}$ given above, the smallest Josephson junctions that we could fabricate while ensuring symmetry of the SQUID at the $\sim 1~{}\%$ level would yield a sufficiently large four-to-one photon exchange rate to enable inflation of a coherent Schrödinger cat state from vacuum [10]). We therefore targeted this minimum value of $E_{J}=h\times$5 GHz.
@@
${\varphi_{a}^{ZPF}}$ 0.45 0.405
$\varphi_{b}^{ZPF}$ 0.55 0.312
$\varphi_{r}^{ZPF}$ 0.05 0.044
@@
Table 1: Targeted and measured system parameters. Note that the estimated Josephson energy in our experiment is three times larger than its target value, entailing strong non-linearity of all modes (see Sec. S2). The achieved dissipation rate for the buffer is also significantly smaller than its nominal values, which limits the engineered four-photon dissipation rate (see Fig. 4).
S1.2 Fabrication process
Wafer preparation: The circuit is fabricated from a 430 $\mu$m-thick wafer of 0001-oriented, double-side epipolished Sapphire C. The sapphire wafer is initially cleaned through a stripping process in a reactive ion etching (RIE) machine, after which it is loaded into a sputtering system. After one night of pumping, we initiate an argon milling cleaning step, followed by the sputtering of 120 nm of niobium. Subsequently, we apply a protective layer of optical resist (S1805), dice the wafer and clean the small chips in solvents.
@@
Aluminum deposition: The chip is then loaded in an e-beam evaporator. We start with a thorough argon ion milling for 2 min at ±$\pm 30^{\circ}$ angles. We then evaporate 35 nm and 70 nm of aluminum, at ±$\pm 30^{\circ}$ angles, separated by an oxidation step in 100 mbar of pure oxygen for 10 min. After evaporating the two layers, the sample is oxidized for 5 min in 100 mbar of oxygen to provide a controlled oxide cap.
Sample mounting: The chip was subsequently glued with PMMA onto a PCB, wire-bonded and mounted into a sample holder. The device was then thermally anchored to the base plate of a Bluefors dilution refrigerator, surrounded by three concentric cans for magnetic and infrared shielding (outer: cryoperm, middle: aluminum, inner: copper, see Fig. S3).
@@
S1.3 Circuit layout details and measurement setup
Figure S1: Optical microscope pictures of the circuit (false colors). The three-pad structure supporting the memory and buffer modes is visible in blue and orange in the zoomed-in picture on the left (corresponding to the purple rectangle in the zoomed-out picture on the right). The readout resonator is colored in green. The width of the center track of the charge line (light blue) and that of the flux lines (mauve) is sinusoidally modulated to prevent radiative decay of the memory (see Fig. S2). The lengths of the flux lines are not exactly matched.
@@
An optical microscope image of our circuit is shown in Fig. S1. The core part of the circuit (zoomed-in picture), comprising the three-pad structure hosting the memory and buffer, the ATS and the termination of the feedlines and readout resonator, is made out of evaporated aluminum. The rest of the circuit (zoomed-out picture), comprising the ground plane, the feedlines and the readout resonator is made out of sputtered niobium. Most of the chip is covered by long meandering coplanar waveguide (CPW) transmission lines. The width of the center track of the lines is periodically modulated to limit radiative decay of the memory, which we now detail.
@@
A 1D microwave transmission line is characterized by its inductance $l$ and capacitance $c$ per unit length. If these electrical properties are periodically modulated with a spatial period $2\pi/\eta$, one obtains a microwave photonic crystal featuring a gap in its energy spectrum. More specifically, for a small modulation amplitude $\epsilon\ll 1$ of both the inductance per unit length $l(x)=l(1+\epsilon\cos(\eta x))$ and the capacitance per unit length $c(x)=c(1+\epsilon\cos(\eta x))^{-1}$—yielding a constant wave propagation velocity $v=1/\sqrt{lc}$ and a modulated characteristic impedance $Z(x)=Z(1+\epsilon\cos(\eta x))$—the gap is centered at $\omega_{G}=v\eta/2$ and its width increases with $\epsilon$ [77]  [^9] This formula holds in the more general case where $v$ also varies, albeit after replacing $v$ with its average value.. A consequence of this gapped spectrum is that the line behaves as a stopband filter for incoming microwave signals, whose in-gap isolation increases exponentially with the number of modulation periods. Note that Stepped impedance filters [79], which are the analog of Bragg mirrors used in optics, were previously employed to this same end [80]. They may be seen as photonic crystals with multiple spatial frequencies $(2n+1)\eta$ such that stopbands appear in their transmission at $(2n+1)v\eta/2$. Photonic crystals were also employed in traveling-wave parametric amplifiers to enable phase-matching [81].
@@
Figure S2: Photonic crystal filters. a) Schematic representation of the charge line CPW geometry, featuring the width of its center track and gaps. Metalized areas are colored in gray (the ground planes at the top and bottom extend beyond the graphic frame, and are connected together by wirebonds). The flux lines layout is identical, except for a reduced length (two periods of modulation instead of five). b) Estimated characteristic impedance of the line, which is locally set by the center track width. c) Numerically computed transmission coefficient of the charge line when connected to 50 $\Omega$ ports (blue). The stopband center frequency is set by the line modulation period, and its width by the amplitude of modulation of its impedance. The flux lines, which are shorter, are not as strongly attenuated in-band.
In our sample, all three feedlines are filtered with these photonic crystals. As illustrated in Fig. S2a, the modulation period is $142\pi/\eta=$ 14 mm for a stopband centered at 4.5 GHz (expected memory frequency). The total CPW width is held constant at 200 $\mu$m, and the inner conductor width is sinusoidally varied from $17~{}\mu$m to $195~{}\mu$m. With these parameters, the line impedance is modulated in the range $20-100~{}\Omega$, and is $50~{}\Omega$ after an integer number of periods. The latter property ensures minimal reflections at the far end of the lines (connected to 50 $\Omega$ ports) for out-of-band microwaves. Note that the impedance variation is not small as in the simple picture presented above, and the modulation is not strictly sinusoidal. Owing to the properties of CPW lines, the phase velocity does not depend significantly on the line geometry.
@@
We perform numerical simulations in order to quantitatively estimate the in-band attenuation of the feedlines. In a nutshell, we use a discrete-cell model for the line (cell electrical length much shorter than relevant wavelengths) and compute the transfer matrix for each cell. We then multiply these matrices to obtain the line transfer matrix. We finally translate the result in a scattering matrix assuming that the ports at each end of the line have an impedance of $50~{}\Omega$ [79]. Intuitively, the transmission coefficient so obtained informs us on the extra protection provided by the modulated line as compared to the case where the system is directly connected to a 50 $\Omega$ port. We report the line transmission against probe frequency in Fig. S2c. The charge line, which is more strongly coupled to the circuit modes, is modulated over 5 periods, yielding more than 40 dB of attenuation. For lack of space on the chip, the flux lines are shorter, with only 2 periods of modulation each, yielding about 15 dB of attenuation. We estimate with electromagnetic simulations that these attenuations are sufficient for radiative decay through the lines not to limit the lifetime of the memory. For all lines, the attenuation varies smoothly over the stopband, whose width exceeds 1 GHz. This feature is particularly appealing as filtering is robust against fabrication disorder, which can result in imprecise targeting of the modes frequency (see Table. 1). Note that the buffer, nominally out of band of the filters, is measured to be close to the stopband. This may partly explain the unexpectedly low value of its single-photon dissipation rate (five times smaller than its nominal value).
@@
Figure S3: Wiring diagram Microwave pulses are generated and detected at room temperature by mixing local oscillators (APUAS from Anapico) with waveforms in the $10-200$ MHz range generated by an OPX from Quantum machines (ADC and DAC ports labeled by their function). DC currents biasing the circuit and SNAIL-parametric amplifier [82] (SPA) are provided by voltage sources from Yokogawa and are thermalized with Therma uD RC filters from Aivon anchored on the 4 K plate of the dilution cryostat. Microwave input lines are attenuated and filtered, in particular with infrared filters (IR) from Bluefors. Output signals are amplified by the SPA and HEMT amplifiers from Low-noise factory and other amplifiers at room temperature.
S2 Non-idealities and non-linearity of the modes

### v3

@@ Acknowledgments
Appendix A Circuit design and measurement setup
A.1 Target parameters and circuit design
@@
${\bf H}_{4}=\hbar g_{4}({\bf a}^{4}-\alpha^{4}){\bf b^{\dagger}}+\mathrm{h.c.}$ (6)
@@
$\dot{\boldsymbol{\rho}}=\gamma_{4}\mathcal{D}[{\bf a}^{4}-\alpha^{4}]\boldsymbol{\rho}$ (7)
with $\gamma_{4}=4g_{4}^{2}/\kappa_{b}$, and the coherent states spanning the cat manifold are attractors of the dynamics. Note that this property remains true in the regime $g_{4}\gtrsim\kappa_{b}$, even though the two-mode dynamics is no longer accurately captured by (7[58]. In Eq.(6), the term $\alpha^{4}{\bf b^{\dagger}}+\mathrm{h.c.}$ is trivially implemented with a resonant drive on the buffer, so that we now focus on the engineering of the more challenging four-to-one photon exchange interaction.
@@
$\begin{split}{\bf H}(t)=&~\hbar\omega_{a}{\bf a^{\dagger}}{\bf a}+\hbar\omega_{b}{\bf b}{\bf b^{\dagger}}\\ &~+2E_{J}\epsilon(t)\sin\big(\varphi^{ZPF}_{a}({\bf a}+{\bf a^{\dagger}})+\varphi^{ZPF}_{b}({\bf b}+{\bf b^{\dagger}})\big)\end{split}$ (8)
where we have neglected the weakly participating readout mode and considered an ideal circuit for simplicity (see Sec. B for an analysis of non-idealities). Note that this expression is valid in the limit $|\epsilon(t)|\ll 1$ as it results from a linearization of $\sin(\epsilon(t))$ [29].
@@
$\begin{split}e^{i\varphi^{ZPF}_{a}({\bf a}e^{-i\omega_{a}t}+{\bf a^{\dagger}}e^{+i\omega_{a}t})}&=e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}}\sum_{n,m}\frac{(i{\varphi_{a}^{ZPF}})^{n+m}}{n!m!}e^{i(n-m)\omega t}{\bf a^{\dagger}}^{n}{\bf a}^{m}\\ e^{-i{\varphi_{a}^{ZPF}}({\bf a}e^{-i\omega_{a}t}+{\bf a^{\dagger}}e^{+i\omega_{a}t})}&=e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}}\sum_{n,m}\frac{(-i{\varphi_{a}^{ZPF}})^{n+m}}{n!m!}e^{i(n-m)\omega t}{\bf a^{\dagger}}^{n}{\bf a}^{m}\end{split}$ (9)
@@
$\tilde{{\bf H}}_{4}=\hbar g_{4}:P_{a}({\bf a^{\dagger}}{\bf a})P_{b}({\bf b^{\dagger}}{\bf b})~{\bf a}^{4}{\bf b^{\dagger}}:~+~\mathrm{h.c.}$ (10)
where :_::\_: denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi{\varphi_{a}^{ZPF}}^{4}{\varphi_{b}^{ZPF}}e^{-\frac{{\varphi_{a}^{ZPF}}^{2}}{2}-\frac{{\varphi_{b}^{ZPF}}^{2}}{2}}$ and
$\begin{split}P_{a}(X)=4!\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{a}^{ZPF}}^{2n}}{n!(n+4)!}X^{n}\\ P_{b}(X)=\sum_{n\geqslant 0}(-1)^{n}\frac{{\varphi_{b}^{ZPF}}^{2n}}{n!(n+1)!}X^{n}\end{split}$ (11)
Compared to the target Hamiltonian (6), higher-order terms appear (corresponding to terms indexed by $n\geqslant 1$ in $P_{a}$ and $P_{b}$). These terms renormalize the four-to-one photon exchange rate between states $|k+4\rangle_{a}|l\rangle_{b}$ and $|k\rangle_{a}|l+1\rangle_{b}$ for $(k,l)\neq(0,0)$. Note that in our experiment, the modes are non-linear so that the pump is not resonant with any of these transitions. Therefore, the system dynamics is the same as that yielded by (6) (see Sec. B and Sec. D.2).
The four-to-one photon exchange rate $g_{4}$ in (10) depends strongly on $\varphi_{a}^{ZPF}$}, which should thus be maximized. However, when stabilizing a four-legged cat in a linear system, high-order terms in the sine expansion can have a a significant impact on the dynamics when $\varphi^{ZPF}_{a}\alpha\gtrsim 1$ such that the sine potential cannot be truncated at leading order. In practice, we saw in numerical simulations that beyond this limit, the cat blobs are distorted and eventually become unstable (not shown). In particular, we found that one should limit ${\varphi_{a}^{ZPF}}\leq 0.5$ to stabilize a cat containing up to 10 photons on average. As for high-order terms in $P_{b}$, they should have no impact on the system dynamics when it is close to its steady state [63]. However, the prefactor $e^{-\frac{{\varphi_{b}^{ZPF}}^{2}}{2}}$ limits the four-to-one photon exchange rate which reaches a maximum for $\varphi^{ZPF}_{b}=1$. In practice, designing a circuit with modes whose vacuum fluctuations across the ATS approach 1 is challenging and one gains only marginally beyond $\varphi^{ZPF}_{b}\sim 0.5$, so that we targeted in design ${\varphi_{a}^{ZPF}}=0.45$ and ${\varphi_{b}^{ZPF}}=0.55$.
As for the readout mode which we have neglected so far, we targeted ${\varphi_{r}^{ZPF}}=0.05$ which was deemed sufficient to perform photon-number resolved measurements of the memory leveraging the non-linearity of the ATS [82] while ensuring that the memory anharmonicity remained smaller than its linewidth at all the ATS DC bias points. This last property is convenient to characterize the system by monitoring the readout mode frequency as a function of bias fluxes (see Fig. 2).
@@
The mode resonance frequencies are not particularly constrained toward the goal of stabilizing a four-legged cat. However, the resonance frequencies should be sufficiently separated that the feedlines can be filtered to obtain a strong single-photon dissipation on the buffer while retaining a large intrinsic quality factor for the memory. Moreover, we generate the pump microwave with standard microwave equipment (in the 1201-20 GHz range) and should avoid frequency collisions with processes resonant at $n\omega_{a}+m\omega_{b}$ with $(n,m)\neq(4,-1)$. In practice, we targeted $\omega_{a}=2\pi\times 4.5$~GHz and $\omega_{b}=2\pi\times 8$~GHz. For the resonance frequency of the readout mode, we aimed at the center of the working range of our parametric amplifier (see Fig. 7).
@@
The four-to-one photon exchange rate $g_{4}$ scales linearly with $E_{J}$ so that naively, one would aim for large $E_{J}$ values to increase the engineered dissipation rate. However, spurious non-linearities stemming from the tunneling of pairs of Cooper pairs through the SQUID scales with $E_{J}^{2}$ (see Sec. B). Through numerical simulations, we estimated that for the target values of $\varphi_{a}^{ZPF}$} and $\varphi_{b}^{ZPF}$} given above, the smallest Josephson junctions that we could fabricate while ensuring symmetry of the SQUID at the $\sim 1~\%$ level would yield a sufficiently large four-to-one photon exchange rate to enable inflation of a coherent Schrödinger cat state from vacuum [36]). We therefore targeted this minimum value of $E_{J}=h\times$5 GHz.
@@
$\varphi_{a}^{ZPF}$} (SQUID) 0.43 0.405
$\varphi_{b}^{ZPF}$ (SQUID) 0.51 0.312
$\varphi_{r}^{ZPF}$ (SQUID) 0.05 0.044
$\varphi_{a,\mathrm{series}}^{ZPF}$} (series inductor) 0.0012 0.0037
$\varphi_{b,\mathrm{series}}^{ZPF}$ (series inductor) 0.0014 0.0028
$\varphi_{r,\mathrm{series}}^{ZPF}$ (series inductor) 0.00014 0.0004
@@
Table 2: Targeted and measured system parameters. The phase vacuum fluctuations of each mode are given across the whole SQUID and across the inductors in series with the Josephson junctions (see Fig. 2a). These inductors account for the inductance of the SQUID arms in our model. As detailed in Sec. B, their inductance is either estimated from electromagnetic simulations (for the target values, estimated inductance 80 pH) or from a fit of the memory anharmonicity pattern (measured value, estimated inductance 85 pH). Note that the estimated Josephson energy in our experiment is three times larger than its target value, which explains that a larger fraction of the modes’ phase fluctuations is allocated to the series inductors for the measured values than for the target values. This stronger participation of the series inductors entails a stronger non-linearity of all modes  than anticipated (see Sec. B). The achieved dissipation rate for the buffer is also significantly smaller than its nominal values, which limits the engineered four-photon dissipation rate (see Fig. 4).
A.2 Fabrication process
[…]
@@ A.2 Fabrication process
Aluminum deposition: The chip is then loaded in an e-beam evaporator. We start with a thorough argon ion milling for 2 min at $\pm 30^{\circ}$ angles. We then evaporate 35 nm and 70 nm of aluminum, at $\pm 30^{\circ}$ angles, separated by an oxidation step in 100 mbar of pure oxygen for 10 min. After evaporating the two layers, the sample is oxidized for 5 min in 100 mbar of oxygen to provide a controlled oxide cap.
Sample mounting: The chip was subsequently glued with PMMA onto a PCB, wire-bonded and mounted into a sample holder. The device was then thermally anchored to the base plate of a Bluefors dilution refrigerator, surrounded by three concentric cans for magnetic and infrared shielding (outer: cryoperm, middle: aluminum, inner: copper, see Fig. 7).
@@
A.3 Circuit layout details and measurement setup
Figure 5: Optical microscope pictures of the circuit (false colors). The three-pad structure supporting the memory and buffer modes is visible in blue and orange in the zoomed-in picture on the left (corresponding to the purple rectangle in the zoomed-out picture on the right). The readout resonator is colored in green. The width of the center track of the charge line (light blue) and that of the flux lines (mauve) is sinusoidally modulated to prevent radiative decay of the memory (see Fig. 6). The lengths of the flux lines are not exactly matched.
@@ A.3 Circuit layout details and measurement setup
An optical microscope image of our circuit is shown in Fig. 5. The core part of the circuit (zoomed-in picture), comprising the three-pad structure hosting the memory and buffer, the ATS and the termination of the feedlines and readout resonator, is made out of evaporated aluminum. The rest of the circuit (zoomed-out picture), comprising the ground plane, the feedlines and the readout resonator is made out of sputtered niobium. Most of the chip is covered by long meandering coplanar waveguide (CPW) transmission lines. The width of the center track of the lines is periodically modulated to limit radiative decay of the memory, which we now detail.
@@
A 1D microwave transmission line is characterized by its inductance $l$ and capacitance $c$ per unit length. If these electrical properties are periodically modulated with a spatial period $2\pi/\eta$, one obtains a microwave photonic crystal featuring a gap in its energy spectrum. More specifically, for a small modulation amplitude $\epsilon\ll 1$ of both the inductance per unit length $l(x)=l(1+\epsilon\cos(\eta x))$ and the capacitance per unit length $c(x)=c(1+\epsilon\cos(\eta x))^{-1}$—yielding a constant wave propagation velocity $v=1/\sqrt{lc}$ and a modulated characteristic impedance $Z(x)=Z(1+\epsilon\cos(\eta x))$—the gap is centered at $\omega_{G}=v\eta/2$ and its width increases with $\epsilon$ [68, 73]. A consequence of this gapped spectrum is that the line behaves as a stopband filter for incoming microwave signals, whose in-gap isolation increases exponentially with the number of modulation periods. Note that Stepped impedance filters [50], which are the analog of Bragg mirrors used in optics, were previously employed to this same end [6]. They may be seen as photonic crystals with multiple spatial frequencies $(2n+1)\eta$ such that stopbands appear in their transmission at $(2n+1)v\eta/2$. Photonic crystals were also employed in traveling-wave parametric amplifiers to enable phase-matching [49].
@@
Figure 6: Photonic crystal filters. a) Schematic representation of the charge line CPW geometry, featuring the width of its center track and gaps. Metalized areas are colored in gray (the ground planes at the top and bottom extend beyond the graphic frame, and are connected together by wirebonds). The flux lines layout is identical, except for a reduced length (two periods of modulation instead of five). b) Estimated characteristic impedance of the line, which is locally set by the center track width. c) Numerically computed transmission coefficient of the charge line when connected to 50 $\Omega$ ports (blue). The stopband center frequency is set by the line modulation period, and its width by the amplitude of modulation of its impedance. The flux lines, which are shorter, are not as strongly attenuated in-band.
[…]
@@
We perform numerical simulations in order to quantitatively estimate the in-band attenuation of the feedlines. In a nutshell, we use a discrete-cell model for the line (cell electrical length much shorter than relevant wavelengths) and compute the transfer matrix for each cell. We then multiply these matrices to obtain the line transfer matrix. We finally translate the result in a scattering matrix assuming that the ports at each end of the line have an impedance of $50~\Omega$ [50]. Intuitively, the transmission coefficient so obtained informs us on the extra protection provided by the modulated line as compared to the case where the system is directly connected to a 50 $\Omega$ port. We report the line transmission against probe frequency in Fig. 6c. The charge line, which is more strongly coupled to the circuit modes, is modulated over 5 periods, yielding more than 40 dB of attenuation. For lack of space on the chip, the flux lines are shorter, with only 2 periods of modulation each, yielding about 15 dB of attenuation. We estimate with electromagnetic simulations that these attenuations are sufficient for radiative decay through the lines not to limit the lifetime of the memory. For all lines, the attenuation varies smoothly over the stopband, whose width exceeds 1 GHz. This feature is particularly appealing as filtering is robust against fabrication disorder, which can result in imprecise targeting of the modes frequency (see Table. 2). Note that the buffer, nominally out of band of the filters, is measured to be close to the stopband. This may partly explain the unexpectedly low value of its single-photon dissipation rate (five times smaller than its nominal value).
@@
Figure 7: Wiring diagram Microwave pulses are generated and detected at room temperature by mixing local oscillators (APUAS from Anapico) with waveforms in the 1020010-200 MHz range generated by an OPX from Quantum machines (ADC and DAC ports labeled by their function). DC currents biasing the circuit and SNAIL-parametric amplifier [13] (SPA) are provided by voltage sources from Yokogawa and are thermalized with Therma uD RC filters from Aivon anchored on the 4 K plate of the dilution cryostat. Microwave input lines are attenuated and filtered, in particular with infrared filters (IR) from Bluefors. Output signals are amplified by the SPA and HEMT amplifiers from Low-noise factory and other amplifiers at room temperature.
Appendix B Non-idealities and non-linearity of the modes